OpenStudy (anonymous):

What are the excluded values? 2/x(x-3)

OpenStudy (anonymous):

here are options x = 0 and 3 x = 0 and 2 x = -3 and 3 none of the above

OpenStudy (anonymous):

Thevalues where the denominator = 0 are at x = 0 and x = 3. So these values must be excluded.

OpenStudy (anonymous):

any more?

OpenStudy (anonymous):

thx what about this one.

OpenStudy (anonymous):

/Users/nchoglo/Desktop/Screen Shot 2014-10-29 at 4.10.08 PM.png

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

poped up

OpenStudy (anonymous):

Identify the polynomial. a ^2 b - cd^3 a.monomial b.binomial c.trinomial d.four-term polynomial e.five-term polynomial

OpenStudy (anonymous):

c.trinomial

OpenStudy (anonymous):

thx what about Find the sum of the given polynomials. a - b + c, b - c + d, and c - d + e a + c + e a + 2b + 3c + 2d + e b + d + e a - c - e

OpenStudy (anonymous):

a + 2b + 3c + 2d + e

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

what about Take 4x + 2 from 8x + 5. -4x - 3 4x + 3 4x - 3

OpenStudy (anonymous):

4x2 - 8x - 5 = 4x2 + 2x -10x -5 = 2x(2x+1) - 5(2x+1) = (2x+1)(2x-5)

OpenStudy (anonymous):

what about What is the middle product of (-2x + 1)(-3x - 5)? 7x -3x -7x

OpenStudy (anonymous):

-7x

OpenStudy (anonymous):

thank you vary much thats it :)

OpenStudy (anonymous):

ok